# Born-Infeld Electrogravity and Dyonic Black Holes

### v1

@@ Abstract
Born-Infeld electrogravity is defined through a Lagrangian that couples gravity and electromagnetism within a single determinantal structure. The field equations are derived in Palatini’s formalism, where the metric, connection, and vector potential are varied independently in the action. As a result, the gravitational sector reduces to Einstein’s equations with a torsion-free, metric-compatible connection. The electrodynamic sector, in turn, admits two equivalent interpretations or pictures: it can be seen either as a standard Born-Infeld electrodynamics in an effective background geometry, or as an anomalous Born-Infeld electrodynamics in the physical metric. We illustrate the dynamics by analyzing the horizon structure and extremality conditions of spherically symmetric dyonic solutions. A comparison with the Reissner-Nordström geometry shows that Born-Infeld electrogravity softens but does not eliminate curvature divergences, and geodesic incompleteness persists.
@@ I Introduction
Born-Infeld theory (BI) has its origins in the context of nonlinear electrodynamics. Proposed by Max Born and Leopold Infeld in 1934 [1, 2, 3, 4], this theory emerged as an extension of classical Maxwell electromagnetism, with the aim of removing the divergences in the self-energy of the point-like charged particle by the following action
$S\propto\int d^{4}x\sqrt{-\det\left(g_{\mu\nu}+bF_{\mu\nu}\right)}-\lambda\sqrt{-\det(g_{\mu\nu})},$ (1)
@@
$S\propto\int d^{4}x\sqrt{-\det\left(g_{\mu\nu}+bR_{\mu\nu}\right)}-\lambda\sqrt{-\det(g_{\mu\nu})}.$ (2)
@@
where $M_{\mu\nu}$ denotes a tensor constructed from matter fields, and $\epsilon$ and $\beta$ (the inverse of the Born-Infeld constant $b$) are coupling constants. For massive scalar fields one may take $M_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi$, while for electromagnetic fields $M_{\mu\nu}=F_{\mu\nu}$. A construction of this type does not fit into the usual separation of matter and gravity in theories based on the principle of minimal coupling, which suggests that such models should eventually conflict with the experimental evidence supporting the equivalence principle. Nevertheless, Vollick showed that for $M_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi$, the field equations coincide with those of GR minimally coupled to a free scalar field. For electromagnetic and fermionic fields, at the first non-trivial order in perturbations, one recovers the usual Einstein-Maxwell and Einstein-Dirac systems, indicating that such theories effectively reduce to an MTG representation at that order.
The main purpose of this work is to investigate whether Born-Infeld electrogravity corrections can smooth or avoid the inherent singularities of GR solutions, while keeping the asymptotic behavior unchanged. For that purpose, we employ the gravito-electromagnetic Lagrangian by following the pioneering Vollick proposal (Eq. (3)), recently revisited by Afonso et al [15], with $M_{\mu\nu}=F_{\mu\nu}$ and we deduce the field equation by varying the action à la Palatini (see, for instance, [16]), where the connection and the metric are treated as independent variables. In particular, we analyze here a dyonic black hole, a spherically symmetric spacetime with both electric and magnetic charges, and find that the singularity is partially smoothed.
The dynamical equations for the gravitational field reduce to Einstein’s equations with a torsion-free, metric-compatible connection. In turn, the electromagnetic field equations admit two equivalent interpretations or pictures: (i) a standard Born-Infeld electrodynamics in an effective background geometry, or (ii) an anomalous Born-Infeld electrodynamics in the physical metric. This BI electrogravity ends up being equivalent to General Relativity sourced by the anomalous Born-Infeld field. Suggestively, this resembles the correspondence between modified gravity and General Relativity plus different matter fields [17, 18]. To the best of our knowledge, the formalism for relating both pictures is presented here for the first time.
The article is organized as follows. Section II introduces the gravito-electromagnetic Lagrangian with non-minimal coupling (Eq. (3)), and we obtain the field equations in the Palatini formalism. Section III exploits the structure of Born–Infeld-like Lagrangians to obtain the dynamics. Section IV shows that, despite appearances, the system evolves consistently with an MTG theory. Section V exhibits the results concerning the dyonic black hole, which carries both electric and magnetic charges. The horizons and singularities are analyzed; in the asymptotic weak field regime, the solution is shown to approach the Reissner–Nordström geometry. Finally, the conclusions are presented in Section VI.
II Born-Infeld electrogravity
II.1 Born-Infeld electrodynamics
@@
Born-Infeld Lagrangian was initially introduced as an idea to avoid the divergence of the electrostatic field of a point-like charge [1, 2, 3]. In its most primitive version, it simply mimicked the Lagrangian of relativistic mechanics, since its electrostatic form was
@@
This was a successful attempt, in the sense that the electric field $\mathbf{E}$ of a point-like charge turned out to be bounded above by the new Born-Infeld constant $b$, in the same way that $c$ is an upper limit for speed in relativistic mechanics. Born and Infeld soon realized that this Lagrangian was the Minkowskian version of a more general one, whose geometric formulation in terms of the electromagnetic field $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is given by the density [4]
@@
Born-Infeld electrodynamics can be summarized through its dynamical equation and energy-momentum tensor, [19, 20, 21, 22]
@@
II.2 Born-Infeld electrogravity
Inspired by Born-Infeld nonlinear electrodynamics, and motivated by its capabilities of smoothing singularities, we will study an electrogravity theory governed by the Lagrangian [14, 15]
@@ II.2 Born-Infeld electrogravity
The parameters $\epsilon$ and $\beta$ are fundamental constants of the theory, with dimensions of squared length and inverse field strength, respectively. $\epsilon^{-1}\beta^{2}$ is proportional to the Newton constant; $\lambda\neq 1$ indicates the presence of a cosmological constant $\Lambda=\epsilon^{-1}(1-\lambda)$, as will be shown.
@@ II.3 Dynamical equations à la Palatini
In the Palatini formalism [16], $\bf g$}, $\bf\Gamma$}, and $\bf A$} are independent dynamical variables. The variation of the Lagrangian with respect to each of them leads to three equations governing the coupled dynamics of the metric, the affine connection, and the electromagnetic field.
@@
So, the symmetric and antisymmetric parts of the tensor density $\sqrt{-q}\;\bar{q}^{\mu\nu}$ separately enter the constraint equation (16) and the dynamical equation (17), as a nice consequence of the symmetric and antisymmetric characters of $\delta g_{\mu\nu}$ and $\delta F_{\mu\nu}$ respectively.
@@
Notice that the symmetric character attributed to $R_{\mu\nu}$ forces the variation $\delta{\bf\Gamma}$ to ensure that $\delta R_{[\mu\nu]}=0$; therefore, the variation $\delta{\bf\Gamma}$ cannot be considered arbitrary. However, the projective-type contribution [23, 24] to $\delta R_{[\mu\nu]}$ coming from $\delta\Gamma^{\alpha}_{\mu\alpha}$ in the second term of Eq. (18) should not worry; its contribution to the variation of the Lagrangian is degenerate with the one coming from $\delta A_{\mu}$ and can be absorbed (see Appendix B). Actually, condition $\delta R_{[\mu\nu]}=0$ can be enforced by just fixing the torsion zero; in this way it is $\delta\Gamma^{\alpha}_{[\mu\nu]}=0$ (because the torsion is fixed), and no contributions to $\delta R_{[\mu\nu]}$ from the first and third terms in Eq. (18) will apprear. So Eq. (18) becomes
@@
and the corresponding variation of the Lagrangian yields the dynamical equation  [^3] As follows from the integration by parts leading to this equation, $\nabla^{\Gamma}_{\alpha}\left(\sqrt{-q}\,\xi^{\nu\mu}\right)$ denotes the tensor density $\sqrt{-q}\,\nabla^{\Gamma}_{\alpha}\xi^{\nu\mu}+\xi^{\nu\mu}\,\partial_{\alpha}\sqrt{-q}$.
@@
However, the constraint (16) turns this equation into the condition for the connection being metric-compatible. Combined with the torsion-free condition imposed above, this uniquely determines the connection to be the Levi-Civita connection. Thus, Eqs. (16) become dynamic, since it now contains second derivatives of the metric. Equations (16) and (17) govern the dynamics of the metric and electromagnetic fields. To make practical use of this system of equations, one must either explicitly construct the tensor $\bar{q}^{\mu\nu}$, or find a shortcut to avoid that task. In this sense, we may notice that from the electrodynamical point of view, the Lagrangian (9)–(10) corresponds to a Born-Infeld electrodynamics in a geometry where the symmetric tensor
@@
plays the role of the background metric. This means that Eq. (17) must correspond to a Born-Infeld equation in the “metric” $\mathcal{G}_{\mu\nu}$. On the other hand, the variation of the Lagrangian (9)–(10) with respect to the metric $g_{\mu\nu}$ should be related to the energy-momentum tensor of the Born-Infeld field in the background geometry $\mathcal{G}_{\mu\nu}$. Since Born-Infeld electrodynamics in arbitrary geometric backgrounds is well known, these elements could help to gain insight into several aspects of Born-Infeld electrogravity, before undertaking the task of building the tensor $\bar{q}^{\mu\nu}$.
III Dynamics of Born-Infeld electrogravity
@@ III.1 The inverse of $q_{\mu\nu}$
We will end this preparatory subsection by noting that
@@ III.2 Electrodynamics
$\partial_{\mu}\left(\sqrt{-\mathcal{G}}\ \frac{\breve{F}^{\mu\nu}-\beta^{2}\ \breve{P}\ ^{\ast}\breve{F}^{\mu\nu}}{\sqrt{1+2\beta^{2}\ \breve{S}-\beta^{4}\ \breve{P}^{2}}}\right)=0\ .$ (37)
As anticipated, this equation governs the standard Born-Infeld electrodynamics in the geometric background described by the “metric” $\mathcal{G}_{\mu\nu}$, since we recognize in parentheses the respective tensor $\breve{\mathcal{F}}^{\mu\nu}$ as introduced in Eq. (8). Certainly, we do not yet know $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{\mu\nu}$ and its inverse $\bar{\mathcal{G}}^{\mu\nu}$, since they are subject to the gravitodynamic equations.
@@ III.3 Gravitodynamics
$\lambda\,\frac{\delta^{\mu}_{\nu}+\epsilon\,R^{\mu}_{\ \nu}}{\sqrt{\det(\delta^{\mu}_{\nu}+\epsilon\,R^{\mu}_{\ \nu})}}=\frac{\delta^{\mu}_{\nu}+\beta^{2}~{}^{\ast}\breve{F}_{~\lambda}^{\mu}~{}^{\ast}\breve{F}^{\lambda}_{\ \nu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}\ .$ (40)
@@ IV Equivalence principle
Equations (37) and (40) seem to imply that the equivalence principle is violated by the action (9). This impression is caused by the presence of the Ricci tensor $R_{\mu\nu}$ in Eq. (37) which governs the dynamics of the electromagnetic field. In fact, the Ricci tensor plays a role in the volume $\sqrt{-\mathcal{G}}$ and the tensor $\breve{F}^{\mu\nu}=\bar{\mathcal{G}}^{\mu\lambda}\bar{\mathcal{G}}^{\nu\rho}F_{\lambda\rho}$. In addition, the source in the r.h.s of Eq. (40) is also contaminated by the Ricci tensor. However, as shown in Ref. [15], the dynamical equations can be combined in such a way that both undesired contributions of the Ricci tensor disappear. We will use the rest of this subsection to demonstrate this property.
Let us begin with Eq. (38), which is the result of varying the Lagrangian (9) with respect to the metric $g_{\mu\nu}$. This dynamical equation can be regarded as the relation between the operations of raising indices with $\bar{\mathcal{G}}^{\mu\nu}$ or $g^{\mu\nu}$. Although the term $^{\ast}F\ ^{\ast}F$ on the l.h.s of Eq. (38) is annoying, its presence can be easily handled with the help of Eqs. (27) and (36), depending on whether $^{\ast}F\ ^{\ast}F$ acts on $F$ or $^{\ast}F$. For instance, by contracting Eq. (38) with $F_{\nu\rho}$ and using Eq. (27) one obtains
@@
$\lambda^{2}\ \breve{P}=P\ ,\ \ \ \ \ \ \ \ \sqrt{-\mathcal{G}}=\lambda^{2}\sqrt{-g}\ .$ (47)
In this way, the general relation (34) is dynamically converted into
@@
Now, let us focus on a calculation concerning the structure of Born-Infeld field equation (8). Eqs. (48), (43) and (47) can be combined to yield
@@
This equation can be multiplied by volume $\sqrt{-\mathcal{G}}$ (so $\sqrt{-g}$ will appear in the l.h.s., according to Eq. (47)) to get on the right side the density tensor that enters the electrodynamic equation (37).
Equation (52) is a direct consequence of applying the dynamical equation (38). It has a profound meaning; the electrodynamics that emerges from Lagrangian (9) can be understood both as the standard Born-Infeld electrodynamics in the geometric background of “metric” $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{\mu\nu}$ (where the electromagnetic field seems to be coupled to the curvature), or as an anomalous Born-Infeld electrodynamics in the geometric background of metric $g_{\mu\nu}$,
$\partial_{\mu}\left(\sqrt{-{g}}\ \frac{{F}^{\mu\nu}+(\beta/\lambda)^{2}\;P\ ^{\ast}{F}^{\mu\nu}}{\sqrt{1-2(\beta/\lambda)^{2}\;S-(\beta/\lambda)^{4}\;P^{2}}}\right)=0\ .$ (53)
We distinguish these two possibilities by calling them picture $\mathcal{G}$ and picture $g$ respectively. We say that the picture $g$ is anomalous because $S$ appears with the “wrong” sign in the l.h.s. of Eq. (52) . So, for passing from the (standard) r.h.s. to the (anomalous) l.h.s. in Eq. (52) not only the metric must be changed, but $\beta^{2}$ must be replaced with $(\beta/\lambda)^{2}$ (or, alternatively, $\beta^{2}\rightarrow-\beta^{2}$, $F_{\mu\nu}\rightarrow\lambda^{-1}F_{\mu\nu}$). However, the picture $g$ is not intrinsically anomalous, since we could have started from a Lagrangian built with an imaginary value for $\beta$; in this way the anomalous character would be transferred to the picture $\mathcal{G}$, while the l.h.s of Eq. (52) would look as the standard Born-Infeld electrodynamics for the field $\lambda^{-1}F_{\mu\nu}$. Imaginary values for $\beta$ are not forbidden in this theory, since just even powers of $\beta$ appear in the dynamical equations (the Lagrangian (9) is made of even powers of $\beta$ according to Eq. (104) ).
@@
The r.h.s. is a necessarily conserved tensor. But it is not the electromagnetic energy-momentum tensor in picture $\mathcal{G}$ (nor is the l.h.s. of the equation the Einstein tensor of that picture). Therefore, we will try to give the r.h.s. the form of the energy-momentum electromagnetic tensor in the picture $g$. The energy-momentum (7) of standard Born-Infeld electrodynamics can be rewritten, among other equivalent ways, as
@@
So, by applying Eq. (51) it results that
@@
Thus, the source in Eq. (55) is the anomalous energy-momentum tensor of picture $g$:
$G^{\mu}_{\nu}-\Lambda\;\delta^{\mu}_{\nu}=\frac{4\pi\beta^{2}}{\epsilon\lambda}\ {T^{anom}}^{\mu}_{\;\nu}\ .$ (60)
As already stated, Born-Infeld electrogravity depends only on the even powers of $\beta$. In case an imaginary value for $\beta$ is chosen –to avoid the anomalous behavior in the picture gg– the sign of $\epsilon$ must be consistently changed to preserve the positive sign of Newton’s constant in the above equation (cf. [25]). In such a case, the cosmological constant in the former equation must be understood as $(1-\lambda)/(-\epsilon)$.

### v2

# Born–Infeld Electrogravity and Dyonic Black Holes

@@ Abstract
Born–Infeld electrogravity is defined through a Lagrangian that couples gravity and electromagnetism within a single determinantal structure. The field equations are derived in Palatini’s formalism, where the metric, connection, and vector potential are varied independently in the action. As a result, the gravitational sector reduces to Einstein’s equations with a torsion-free, metric-compatible connection. The electrodynamic sector, in turn, admits two equivalent interpretations or pictures: it can be seen either as a standard Born–Infeld electrodynamics in an effective background geometry, or as an anomalous Born–Infeld electrodynamics in the physical metric. We illustrate the dynamics by analyzing the horizon structure, the extremality conditions, and the thermodynamics of spherically symmetric dyonic solutions. Remarkably, in the small-charge limit, Born–Infeld electrogravity admits a fundamental extremal black hole whose mass and horizon area are determined exclusively by the Born–Infeld and Newton constants and the speed of light.
@@ I Introduction
Born–Infeld theory (BI) has its origins in the context of nonlinear electrodynamics. Proposed by Max Born and Leopold Infeld in 1934 [1, 2, 3, 4], this theory emerged as an extension of classical Maxwell electromagnetism, with the aim of removing the divergences in the self-energy of the point-like charged particle by the following action
$S\propto\int d^{4}x\sqrt{-\det\left(g_{\mu\nu}+b^{-1}F_{\mu\nu}\right)}-\lambda\sqrt{-\det(g_{\mu\nu})},$ (1)
@@
$S\propto\int d^{4}x\sqrt{-\det\left(g_{\mu\nu}+\epsilon R_{\mu\nu}\right)}-\lambda\sqrt{-\det(g_{\mu\nu})}.$ (2)
@@
where $M_{\mu\nu}$ denotes a tensor constructed from matter fields, and $\epsilon$ and $\beta$ are universal constants of the theory. For massive scalar fields one may take $M_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi$, while for electromagnetic fields $M_{\mu\nu}=F_{\mu\nu}$. A construction of this type does not fit into the usual separation of matter and gravity in theories based on the principle of minimal coupling, which suggests that such models should eventually conflict with the experimental evidence supporting the equivalence principle. Nevertheless, Vollick showed that for $M_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi$, the field equations coincide with those of GR minimally coupled to a free scalar field. For electromagnetic and fermionic fields, at the first non-trivial order in perturbations, one recovers the usual Einstein-Maxwell and Einstein-Dirac systems, indicating that such theories effectively reduce to an MTG representation at that order.
In this work we follow Vollick’s pioneering proposal of Eq. (3), with $M_{\mu\nu}=F_{\mu\nu}$, recently revisited by Afonso et al [15]. To derive the field equations we vary the action à la Palatini, where the connection and the metric are treated as independent variables. In the standard metric formalism, instead, the connection is given in terms of the metric (namely, the Levi-Civita connection), so that the two objects are not independent. It is well known that Palatini and metric formalisms lead to the same dynamics in General Relativity, Lovelock theories [16] and other particular cases where the dynamical equations end up being second order because the higher derivatives go to boundary terms in the metric formalism. But in general, the metric and Palatini formalisms are two ways of varying the action that lead to different dynamics. In many modified gravity theories, varying the action within the metric formalism leads to fourth-order dynamical equations for the metric, potentially introducing extra degrees of freedom and instabilities such as Ostrogradski ghosts (see [17] and references therein). By contrast, varying the action à la Palatini [18] typically leads to second-order field equations for the metric, thus avoiding the higher-derivative pathologies that appear in the metric approach. However, the Palatini approach is not entirely free of problems. Ghosts may reappear in scalar-tensor theories with matter coupled to the connection (however, see [19, 20]), and the Cauchy problem may be ill-possed [21] (however, see [18]). Nevertheless, these concerns are not relevant for the BI gravito-electromagnetic Lagrangian considered here, since the Palatini variation leads to Einstein equations whose electromagnetic source is a metric energy-momentum tensor. Suggestively, this resembles the correspondence between modified gravity and General Relativity plus different matter fields [22, 23].
The article is organized as follows. Section II introduces the gravito-electromagnetic Lagrangian with non-minimal coupling (Eq. (3)), and we obtain the field equations in the Palatini formalism. Section III exploits the structure of BI-like Lagrangians to obtain the dynamics. Section IV shows that, despite appearances, the system evolves consistently with an MTG theory. We show that the gravitodynamics reduces to Einstein’s equations with a torsion-free, metric-compatible connection, while the electrodynamics admits two equivalent interpretations or pictures: (i) a standard BI electrodynamics in an effective background geometry, or (ii) an anomalous BI electrodynamics in the physical metric. Section V presents the results concerning the dyonic BH, which carries both electric and magnetic charges. We study the horizon structure, the extremality conditions and the thermodynamics. We show that there exist two regimes in the extremal case, depending on the relation between the BH charge and the characteristic charge of the theory, determined by the BI and Newton constants ($\beta$ and $G$) together with the speed of light ($c$). For high charge, the Reissner-Nordström solution is recovered, while for small charges we find a fundamental extremal BH whose mass and horizon area are determined solely by $\beta$, $G$ and $c$. In Section VI, we present the conclusions of the work.
II Born–Infeld electrogravity
II.1 Born–Infeld electrodynamics
@@
Born–Infeld Lagrangian was initially introduced as an idea to avoid the divergence of the electrostatic field of a point-like charge [1, 2, 3]. In its most primitive version, it simply mimicked the Lagrangian of relativistic mechanics, since its electrostatic form was
@@
This was a successful attempt, in the sense that the electric field $\mathbf{E}$ of a point-like charge turned out to be bounded above by the new BI constant $b$, in the same way that $c$ is an upper limit for speed in relativistic mechanics. Born and Infeld soon realized that this Lagrangian was the Minkowskian version of a more general one, whose geometric formulation in terms of the electromagnetic field $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is given by the density [4]
@@
BI electrodynamics can be summarized through its dynamical equation and energy-momentum tensor, [24, 25, 26, 27]
@@
II.2 Born–Infeld electrogravity
Inspired by BI nonlinear electrodynamics, and motivated by its capabilities of smoothing singularities, we study an electrogravity theory governed by the Lagrangian [14, 15]
@@ II.2 Born–Infeld electrogravity
The parameters $\epsilon$ and $\beta$ are fundamental constants of the theory, with dimensions of squared length and inverse field strength, respectively, that satisfy $\epsilon^{-1}\beta^{2}\propto G$, where $G$ is the Newton constant. Finally, $\lambda\neq 1$ indicates the presence of a cosmological constant $\Lambda=\epsilon^{-1}(1-\lambda)$, as will be shown.
@@ II.3 Dynamical equations à la Palatini
Since we adopt the Palatini approach, the variational procedure treats $\bf g$}, $\bf\Gamma$}, and $\bf A$} as independent dynamical variables. Therefore, the variation of the Lagrangian with respect to each of them leads to three equations that govern the coupled dynamics of the metric, the affine connection, and the electromagnetic field.
@@
Thus, the symmetric and antisymmetric parts of the tensor density $\sqrt{-q}\;\bar{q}^{\mu\nu}$ separately enter the constraint equation (16) and the dynamical equation (17), as a nice consequence of the symmetric and antisymmetric characters of $\delta g_{\mu\nu}$ and $\delta F_{\mu\nu}$ respectively.
@@
The symmetric character attributed to $R_{\mu\nu}$ in Eq. (10) forces the variation $\delta{\bf\Gamma}$ to ensure that $\delta R_{[\mu\nu]}=0$; therefore, the variation $\delta{\bf\Gamma}$ is not arbitrary. In Eq. (18), the term containing $\delta\Gamma^{\alpha}_{\mu\alpha}$ effectively contributes to $\delta R_{[\mu\nu]}$; but it is a projective-type contribution [28, 29], which is degenerate with a similar contribution of $\delta A_{\mu}$ to the variation of the Lagrangian (see Appendix B), and can be absorbed into its electromagnetic counterpart. Hence, the condition $\delta R_{[\mu\nu]}=0$ requires only that the torsion be set to zero. In this way, the third term in Eq. (18) does not appear, and $\delta\Gamma^{\alpha}_{[\mu\nu]}$ in the first term vanishes because the torsion has been fixed. [^3] The imposition of zero torsion within the Palatini formalism should not surprise. The proof of the equivalence between the metric and Palatini formalisms in General Relativity and Lovelock theories presupposes that torsion is zero or $\Gamma$ is a metric connection. Thus, Eq. (18) becomes
@@
and the corresponding variation of the Lagrangian yields the dynamical equation  [^4] As follows from the integration by parts that leads to this equation, $\nabla^{\Gamma}_{\alpha}\left(\sqrt{-q}\,\xi^{\nu\mu}\right)$ denotes the tensor density $\sqrt{-q}\,\nabla^{\Gamma}_{\alpha}\xi^{\nu\mu}+\xi^{\nu\mu}\,\partial_{\alpha}\sqrt{-q}$.
@@
However, the constraint (16) turns this equation into the condition for the connection being metric-compatible. Combined with the torsion-free condition imposed above, this uniquely determines the connection to be the Levi-Civita connection. Thus, Eqs. (16) become dynamic, since it now contains second derivatives of the metric. Equations (16) and (17) govern the dynamics of the metric and electromagnetic fields. To make practical use of this system of equations, one must either explicitly construct the tensor $\bar{q}^{\mu\nu}$, or find a shortcut to avoid that task. In this sense, we may notice that from the electrodynamical point of view, the Lagrangian (9)–(10) corresponds to a BI electrodynamics in a geometry where the symmetric tensor
@@
plays the role of the background metric. This means that Eq. (17) must correspond to a BI equation in the “metric” $\mathcal{G}_{\mu\nu}$. On the other hand, the variation of the Lagrangian (9)–(10) with respect to the metric $g_{\mu\nu}$ should be related to the energy-momentum tensor of the BI field in the background geometry $\mathcal{G}_{\mu\nu}$. Since BI electrodynamics in arbitrary geometric backgrounds is well known, these elements could help to gain insight into several aspects of BI electrogravity, before undertaking the task of building the tensor $\bar{q}^{\mu\nu}$.
III Dynamics of Born–Infeld electrogravity
@@ III.1 The inverse of $q_{\mu\nu}$
We end this preparatory subsection by noting that
@@ III.2 Electrodynamics
$\boxed{\partial_{\mu}\left(\sqrt{-\mathcal{G}}\ \frac{\breve{F}^{\mu\nu}-\beta^{2}\ \breve{P}\ ^{\ast}\breve{F}^{\mu\nu}}{\sqrt{1+2\beta^{2}\ \breve{S}-\beta^{4}\ \breve{P}^{2}}}\right)=0\ .}$ (37)
As anticipated, this equation governs the standard BI electrodynamics in the geometric background described by the “metric” $\mathcal{G}_{\mu\nu}$, since we recognize in parentheses the respective tensor $\breve{\mathcal{F}}^{\mu\nu}$ as introduced in Eq. (8). Certainly, we do not yet know $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{\mu\nu}$ and its inverse $\bar{\mathcal{G}}^{\mu\nu}$, since they are subject to the gravitodynamic equations.
@@ III.3 Gravitodynamics
$\boxed{\lambda\,\frac{\delta^{\mu}_{\nu}+\epsilon\,R^{\mu}_{\ \nu}}{\sqrt{\det(\delta^{\mu}_{\nu}+\epsilon\,R^{\mu}_{\ \nu})}}=\frac{\delta^{\mu}_{\nu}+\beta^{2}~{}^{\ast}\breve{F}_{~\lambda}^{\mu}~{}^{\ast}\breve{F}^{\lambda}_{\ \nu}}{\sqrt{1+2\beta^{2}~{S}-\beta^{4}~\breve{P}^{2}}}\ .}$ (40)
@@ IV Equivalence principle
Equations (37) and (40) seem to imply that the equivalence principle is violated by the action (9). This impression is caused by the presence of the Ricci tensor $R_{\mu\nu}$ in Eq. (37) which governs the dynamics of the electromagnetic field. In fact, the Ricci tensor contributes to both the volume $\sqrt{-\mathcal{G}}$ and the tensor $\breve{F}^{\mu\nu}=\bar{\mathcal{G}}^{\mu\lambda}\bar{\mathcal{G}}^{\nu\rho}F_{\lambda\rho}$. In addition, the source in the r.h.s of Eq. (40) is also contaminated by the Ricci tensor. However, as shown in Ref. [15], the dynamical equations can be combined in such a way that both undesired contributions of the Ricci tensor disappear. We will use the rest of this subsection to demonstrate this property.
Let us begin with Eq. (38), which is the result of varying the Lagrangian (9) with respect to the metric $g_{\mu\nu}$. This dynamical equation can be regarded as the relation between the operations of raising indices with $\bar{\mathcal{G}}^{\mu\nu}$ or $g^{\mu\nu}$. Although the term $^{\ast}F\ ^{\ast}F$ on the l.h.s of Eq. (38) is cumbersome, its presence can be easily handled with the help of Eqs. (27) and (36), depending on whether $^{\ast}F\ ^{\ast}F$ acts on $F$ or $^{\ast}F$. For instance, by contracting Eq. (38) with $F_{\nu\rho}$ and using Eq. (27) one obtains
@@
$\lambda^{2}\ \breve{P}=P\ ,\ \ \ \ \ \ \ \ \sqrt{-\mathcal{G}}=\lambda^{2}\sqrt{-g}\ ;$ (47)
the second relationship implies that the “metric” $\mathcal{G}$ is inversible if the metric $g$ is. In this way, the general relation (34) is dynamically converted into
@@
Now, let us focus on a calculation concerning the structure of BI field equation (8). Eqs. (48), (43) and (47) can be combined to yield
@@
This equation can be multiplied by the volume $\sqrt{-\mathcal{G}}$ (so $\sqrt{-g}$ will appear in the l.h.s., according to Eq. (47)) to get on the right side the density tensor that enters the electrodynamic equation (37).
Equation (52) is a direct consequence of applying the dynamical equation (38). It has a profound meaning; the electrodynamics that emerges from Lagrangian (9) can be understood both as the standard BI electrodynamics in the geometric background of “metric” $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{\mu\nu}$ (where the electromagnetic field seems to be coupled to the curvature), or as an anomalous BI electrodynamics in the geometric background of metric $g_{\mu\nu}$,
$\boxed{\partial_{\mu}\left(\sqrt{-{g}}\ \frac{{F}^{\mu\nu}+(\beta/\lambda)^{2}\;P\ ^{\ast}{F}^{\mu\nu}}{\sqrt{1-2(\beta/\lambda)^{2}\;S-(\beta/\lambda)^{4}\;P^{2}}}\right)=0\ .}$ (53)
We distinguish these two possibilities by calling them picture $\mathcal{G}$ and picture $g$ respectively. We say that the picture $g$ is anomalous because $S$ appears with the “wrong” sign in the l.h.s. of Eq. (52) . Therefore, for passing from the (standard) r.h.s. to the (anomalous) l.h.s. in Eq. (52) not only the metric must be changed, but $\beta^{2}$ must be replaced with $(\beta/\lambda)^{2}$ (or, alternatively, $\beta^{2}\rightarrow-\beta^{2}$, $F_{\mu\nu}\rightarrow\lambda^{-1}F_{\mu\nu}$). However, the picture $g$ is not intrinsically anomalous, since we could have started from a Lagrangian built with an imaginary value for $\beta$; in this way the anomalous character would be transferred to the picture $\mathcal{G}$, while the l.h.s of Eq. (52) would look as the standard BI electrodynamics for the field $\lambda^{-1}F_{\mu\nu}$. Imaginary values for $\beta$ are not forbidden in this theory, since only even powers of $\beta$ appear in the dynamical equations (the Lagrangian (9) is made of even powers of $\beta$ according to Eq. (111) ).
@@
The r.h.s. is necessarily a conserved tensor. However, it is not the electromagnetic energy-momentum tensor in picture $\mathcal{G}$ (the l.h.s.either is the Einstein tensor in that picture). Therefore, we will try to give the r.h.s. the form of the energy-momentum electromagnetic tensor in the picture $g$. The energy-momentum (7) of standard BI electrodynamics can be rewritten, among other equivalent forms, as
@@
Therefore, by applying Eq. (51) it results that
@@
Thus, the source in Eq. (55) is the anomalous energy-momentum tensor in picture $g$:
$\boxed{G^{\mu}_{\nu}-\Lambda\;\delta^{\mu}_{\nu}=\frac{4\pi\beta^{2}}{\epsilon\lambda}\ {T^{anom}}^{\mu}_{\;\nu}\ .}$ (60)
As already stated, BI electrogravity depends only on the even powers of $\beta$. In case an imaginary value for $\beta$ is chosen –to avoid the anomalous behavior in the picture gg– the sign of $\epsilon$ must be consistently changed to preserve the positive sign of Newton’s constant in the above equation (cf. [30]). In such a case, the cosmological constant in the former equation must be understood as $(1-\lambda)/(-\epsilon)$.